# Training-free Zero-shot Composed Image Retrieval via Weighted Modality Fusion and Similarity

### v2

@@ Abstract
Composed image retrieval (CIR), which formulates the query as a combination of a reference image and modified text, has emerged as a new form of image search due to its enhanced ability to capture users’ intentions. However, training a CIR model in a supervised manner typically requires labor-intensive collection of (reference image, text modifier, target image) triplets. While existing zero-shot CIR (ZS-CIR) methods eliminate the need for training on specific downstream datasets, they still require additional pretraining with large-scale image-text pairs. In this paper, we introduce a training-free approach for ZS-CIR. Our approach, Weighted Modality fusion and similarity for CIR (WeiMoCIR), operates under the assumption that image and text modalities can be effectively combined using a simple weighted average. This allows the query representation to be constructed directly from the reference image and text modifier. To further enhance retrieval performance, we employ multimodal large language models (MLLMs) to generate image captions for the database images and incorporate these textual captions into the similarity computation by combining them with image information using a weighted average. Our approach is simple, easy to implement, and its effectiveness is validated through experiments on the FashionIQ and CIRR datasets.
@@ 1 Introduction
Multi-modality queries pose a unique challenge to image retrieval because the retrieval algorithms need to understand what to retain in the visual content and what to modify based on finer details expressed in the natural language. Previous studies [3, 13] have tackled this by training CIR models on large-scale (reference image, text modifier, target image) triplets to learn how to fuse the reference image and the text modifier into a query representation for retrieval. However, collecting such triplets requires significant human effort, and these supervised approaches usually exhibit limited performance on unseen data due to being tuned on specific datasets. As a result, zero-shot CIR (ZS-CIR), which is able to perform CIR without the need for supervised learning on downstream data, has gained attention. However, existing ZS-CIR approaches, such as Pic2Word [17] and SEARLE [2], still require training on large image-text datasets to learn mappings for converting images into pseudo-word tokens, which are then combined with textual descriptions for text-to-image retrieval. This additional training can be resource-intensive.
@@ 2 Related Work
In CIR, the query consists of two modalities: a reference image and textual modifications, with goal of retrieving images that meet the conditions specified in the query. Earlier CIR methods [3, 6, 21] have focused on developing multimodal fusion techniques to combine the visual features and textual features to obtain the query representation. For instance, CLIP4CIR [3] trains a combiner to fuse the information from two modalities. However, these methods require triplets for specific downstream applications for model training, and collecting such triplets is labor-intensive. As a result, recent research has explored ZS-CIR. Representative methods, like Pic2Word [17] and SEARLE [2], learn to project images into the textual space using image-text pairs. With the learned mappings from images to texts, the projected image is then combined with the text modifier for text-to-image retrieval without needing to train on downstream datasets. CIReVL [10] further eliminates the need for training on image-text pairs and achieves training-free ZS-CIR. It utilizes a pretrained VLM [11, 15] to generate a textual description for the reference image. This description and the text modifier are fed into an LLM [4, 16] to produce a textual description of the desired image for text-to-image retrieval.

### v3

@@ Abstract
Composed image retrieval (CIR), which formulates the query as a combination of a reference image and modified text, has emerged as a new form of image search due to its enhanced ability to capture users’ intentions. However, training a CIR model in a supervised manner typically requires labor-intensive collection of (reference image, text modifier, target image) triplets. While existing zero-shot CIR (ZS-CIR) methods eliminate the need for training on specific downstream datasets, they still require additional pretraining on large-scale image datasets. In this paper, we introduce a training-free approach for ZS-CIR. Our approach, Weighted Modality fusion and similarity for CIR (WeiMoCIR), operates under the assumption that image and text modalities can be effectively combined using a simple weighted average. This allows the query representation to be constructed directly from the reference image and text modifier. To further enhance retrieval performance, we employ multimodal large language models (MLLMs) to generate image captions for the database images and incorporate these textual captions into the similarity computation by combining them with image information using a weighted average. Our approach is simple, easy to implement, and its effectiveness is validated through experiments on the FashionIQ and CIRR datasets.
@@ 1 Introduction
Multi-modality queries pose a unique challenge to image retrieval because the retrieval algorithms need to understand what to retain in the visual content and what to modify based on finer details expressed in the natural language. Previous studies [3, 13] have tackled this by training CIR models on large-scale (reference image, text modifier, target image) triplets to learn how to fuse the reference image and the text modifier into a query representation for retrieval. However, collecting such triplets requires significant human effort, and these supervised approaches usually exhibit limited performance on unseen data due to being tuned on specific datasets. As a result, zero-shot CIR (ZS-CIR), which is able to perform CIR without the need for supervised learning on downstream data, has gained attention. However, existing ZS-CIR approaches, such as Pic2Word [17] and SEARLE [2], still require training on large image datasets to learn mappings for converting images into pseudo-word tokens, which are then combined with textual descriptions for text-to-image retrieval. This additional training can be resource-intensive.
@@ 2 Related Work
In CIR, the query consists of two modalities: a reference image and textual modifications, with goal of retrieving images that meet the conditions specified in the query. Earlier CIR methods [3, 6, 21] have focused on developing multimodal fusion techniques to combine the visual features and textual features to obtain the query representation. For instance, CLIP4CIR [3] trains a combiner to fuse the information from two modalities. However, these methods require triplets for specific downstream applications for model training, and collecting such triplets is labor-intensive. As a result, recent research has explored ZS-CIR. Representative methods, like Pic2Word [17] and SEARLE [2], learn to project images into the textual space using large-scale image datasets. With the learned mappings from images to texts, the projected image is then combined with the text modifier for text-to-image retrieval without needing to train on downstream datasets. CIReVL [10] further eliminates the need for additional training to learn the image-to-text mappings and achieves training-free ZS-CIR. It utilizes a pretrained VLM [11, 15] to generate a textual description for the reference image. This description and the text modifier are fed into an LLM [4, 16] to produce a textual description of the desired image for text-to-image retrieval.